OpenStudy (anonymous):

The sum of the digits of a two-digit number is 12. If the digits are reversed, the number is decreased by 18. What is the original number?

OpenStudy (anonymous):

75, right?

OpenStudy (anonymous):

I'm not sure.....I can't do it......

OpenStudy (anonymous):

your number XY. x+y = 12 (10x+y) -(10y+x) = 18 just solve this system

OpenStudy (anonymous):

|x and y| <=9

OpenStudy (anonymous):

@bmp is right

OpenStudy (anonymous):

How do you get that? I'm really screwed up.

OpenStudy (anonymous):

your number XY. x+y = 12 (10x+y) -(10y+x) = 18 just solve this system

OpenStudy (anonymous):

the 2 digit number (XY) can be writen like this: 10*x + y = number XY where x y are decimal notation digits form 0...9

OpenStudy (anonymous):

GOT IT

OpenStudy (anonymous):

Just think that every base 10 number is of the form (10^0)*a + (10^1)*b + (10^2)*c... and so on, where 10^0 is the zeroth decimal number, 10^1 is the second etc. So a two digit number is written as 10^0*y + 10^1*x, or 10x + y, as @myko pointed out.

OpenStudy (anonymous):

I understand and I got 75. Thanks!